Given the above demand curve, how many of good X will consumer purchase when PX is $100 a unit, PY is $50 a unit, and M is $25,000?
Answer #1

As it has been given that the above demand curve when PX is $100 a unit, PY is $50 a unit, and M is $25,000.

The equation of the budget constraints are;

Px.X+Py.Y=M

100X+50Y=25,000

50(2X+Y)=25,000

2X+Y=25,000/50

2X=(500-Y)

X=(500-Y)/2 units

Until the quantity purchased of Y good is not given, the quantity of good X cannot be determined. So if the quantity of good Y will be given, then by substituting the value of good Y, the quantity of good X can be determined.